# Level-set shape optimization

### v1

@@ 4.3 PDE-constrained test case
Similarly as in Section 4.2, to solve the level-set equation Eq. 25, we partition the hold-all domain $D$ into 2685 triangles, discretize in space using a classical dG method with basis functions of degree $p=2$, and employ the Heun’s method (with time step $\Delta t_{k}=1/1000$) for time-stepping. Per iteration, we perform at most $M=10$ time steps. To solve the shape gradient equation Eq. 26 and the state equation (28), we agglomerate the locally refined triangular mesh (cf. Fig. 3) into 200 polytopic elements and employ polynomials of degree $p=2$. To assemble the various Galerkin matrices, we employ a quadrature rule of order 4 on triangles, an of order 4 on edges, which computes the involved integrals exactly (up to round-off errors). Note that, as displayed in Fig. 3, the state constraint is solved by extracting its computational domain instead of pursuing the so-called “Ersatz material” approach [2].
Figure 7: Evolution of shapes in numerical experiment 3. The initial domain, a disc, elongates and eventually splits into two domains before converging to the optimal configuration. The contour plots represent the normalized contours of the Euclidean norm of the gradient $\nabla J$. Note that the polytopic mesh changes at each iteration.
@@
$\phi(x)=\phi(x_{1},x_{2})=\sqrt[4]{((x-0.6)^{2}+y^{2})((x+0.6)^{2}+y^{2})}-0.6\,,$
are displayed in Fig. 7. We observe that the algorithm retrieves a good approximation of the optimal shape. To demonstrate the generality of the approach, Fig. 8 displays some shape iterates starting from a different initial guess.
Figure 8: Evolution of shapes in numerical experiment 1 for a different initial domain. The shape iterates happily converge to the optimal annular domain.
@@ 5 Conclusions
Thanks to its flexibility in terms of element shapes, polynomial order, treatment of hanging nodes, stability, the method presented here is perfectly positioned for the development of a posteriori $hp$-adaptive shape optimization methods. Finally, we note that the above developments can be combined with GPU-accelerated 3D implementations of $hp$-version polytopic IPdG methods [19], to enable extremely fast implementation and solution of the RKdG step.

### v2

@@ 4.3 PDE-constrained test case
Similarly as in Section 4.2, to solve the level-set equation Eq. 25, we partition the hold-all domain $D$ into 2085 triangles, discretize in space using a classical dG method with basis functions of degree $p=2$, and employ the Heun’s method (with time step $\Delta t_{k}=1/1000$) for time-stepping. Per iteration, we perform at most $M={100}$ time steps. To solve the shape gradient equation Eq. 26 and the state equation (28), we agglomerate the locally refined triangular mesh (cf. Fig. 3) into 200 polytopic elements and employ polynomials of degree $p=2$. To assemble the various Galerkin matrices, we employ a quadrature rule of order 4 on triangles, an of order 4 on edges, which computes the involved integrals exactly (up to round-off errors). Note that, as displayed in Fig. 3, the state constraint is solved by extracting its computational domain instead of pursuing the so-called “Ersatz material” approach [2].
$n=0$, $t=0$$n=5$, $t=0.5$$n=7$, $t=0.7$$n=11$, $t=1.03$
Figure 8: Evolution of shapes for the PDE-constrained test case. The variables $n$ and $t$ denote the number of gradient evaluations and the pseudo-time, respectively. The two holes in the initial domain merge into one and then expand, converging to the optimal configuration (depicted in red). The contour plots represent the normalized contours of the Euclidean norm of the gradient $\nabla J$. Note that the polytopic mesh changes at each iteration.
@@
$\phi(x)=\phi(x_{1},x_{2})=\sqrt[4]{((x-0.6)^{2}+y^{2})((x+0.6)^{2}+y^{2})}-{0.55}\,,$
are displayed in Fig. 8. We observe that the algorithm retrieves a good approximation of the optimal shape. To demonstrate the generality of the approach, Fig. 9 displays some shape iterates starting from a different initial guess. The convergence histories for both initial guesses are displayed in Fig. 10.
$n=0$, $t=0$$n=5$, $t=0.5$$n=7$, $t=0.7$$n=13$, $t=1.1$
Figure 9: Evolution of shapes for the PDE-constrained test case with a different initial guess. The variables $n$ and $t$ denote the number of gradient evaluations and the pseudo-time, respectively.. The shape iterates happily converge to the optimal annular domain (depicted in red).
Figure 10: Convergence history for the PDE-constrained test case with different initial guesses: one as in Fig. 8 and one as in Fig. 9. The figure on the left depicts the pseudo-time evolution of the difference between the objective value $J$ as shapes are optimized and (an accurate approximation of) the objective value at the optimal solution. The figure on the right depicts the pseudo-time evolution of an approximate distance between the zero-level set of the optimized shapes and the zero-level set of the optimal shape. Markers denote gradient evaluations. We observe that both convergence measures are significantly reduced as optimization proceeds. The convergence lines in the plot on the left plateau due to numerical approximations errors leading to objective values $J$ fictitiously smaller than objective value at the optimal solution.
@@ 5 Conclusions
Thanks to its flexibility in terms of element shapes, polynomial order, treatment of hanging nodes, stability, the method presented here is perfectly positioned for the development of a posteriori $hp$-adaptive shape optimization methods. The above developments can be combined with GPU-accelerated implementations of $hp$-version polytopic IPdG methods [21] to enable extremely fast solution of the RKdG step.
We note that, although the numerical experiments presented are based on two-dimensional shape optimization examples, three-dimensional problems can be treated using the same methodology, without the need of any further considerations. The favorable computational complexity observed is expected to be retained in the context of three-dimensional setting also. Finally, given the availability of interior penalty dG methods for various classes of PDE problems, the treatment of for more complex PDE problems is by all means possible.